# Relaxed bound on performance of quantum key repeaters and secure content of generic private and independent bits

### 29 Feb 2024 v3

@@ IV Relaxed bound on one-way private key repeaters
We follow the proof of the lemma $7$ of Horodecki et al. (2009a). Using concavity of the logarithm function we first note that $D(\psi_{+}^{Rn}||\sigma_{n})=-\operatorname{Tr}\psi_{+}^{Rn}\log\sigma_{n}\geq-\log\operatorname{Tr}\psi_{+}^{Rn}\sigma_{n}=-\log\sum_{i}p_{i}\operatorname{Tr}\psi_{+}^{Rn}\sigma_{n}^{(i)}$. Since $\sigma_{n}\in\mbox{Cone}(SEP,\tilde{\sigma}_{n})$, we have that $\sigma^{(i)}_{n}$ can be considered either separable or equal to $\tilde{\sigma}_{n}$.Suppose $\tilde{\sigma}_{n}$ is admixed in $\sigma_{n}$ with probability $p_{i_{0}}$. For the separable states $\operatorname{Tr}\sigma_{n}^{(i)}\psi_{+}^{Rn}\leq 2^{-Rn}$ by lemma 7 of Horodecki et al. (2009a). Since $Rn\geq n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}E_{\alpha}(\sigma^{\otimes n}))$, we have
$D(\psi_{+}^{Rn}||\tilde{\sigma}_{n})\geq$
$(1-p_{i_{0}})\left[n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}E_{\alpha}(\sigma^{\otimes n}))\right]$
$+p_{i_{0}}\left[-\log\operatorname{Tr}\psi_{+}^{Rn}\tilde{\sigma}_{n}\right]$ (39)
The last step is bounding $-\log\operatorname{Tr}\psi_{+}^{Rn}\tilde{\sigma}_{n}$ by $n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}E_{\alpha}(\sigma^{\otimes n}))$ from below by Lemma 1
We are ready to state the main theorem. It shows that the one-way repeater rate is upper bounded by the locally measured regularized relative entropy of entanglement (increased by factor $\frac{\alpha}{\alpha-1}$) in addition with the sandwich relative entropy of entanglement. In what follows we will use notation $\rho\equiv\rho_{AC_{1}}$ and $\rho^{\prime}\equiv\rho^{\prime}_{C_{2}B}$, as well as $\rho^{\prime\prime}\equiv\rho^{\prime\prime}_{ABC_{1}C_{2}}$ and $C_{1}C_{2}\equiv{\bf C}$.
@@
For any bipartite states $\rho,\rho^{\prime},\rho^{\prime\prime}$ and a real parameter $\alpha\in(1,\infty)$, there is
$R^{\rightarrow}(\rho,\rho^{\prime})\leq\frac{\alpha}{\alpha-1}D_{\bf{C}}^{\infty}(\rho\otimes\rho^{\prime}||\rho^{\prime\prime})+\tilde{E}_{\alpha}(\rho^{\prime\prime}\otimes\rho^{\prime\prime}).$ (40)
@@
Our proof closely follows Christandl and Ferrara (2017), with suitable modifications. We first note, that there is the following chain of (in)equalities, as it is noted in Christandl and Ferrara (2017):
@@
${\frac{1}{n}}D_{\bf{C}}({\cal M}(\rho^{\otimes n}\otimes\rho^{\prime\otimes n})||{\cal M}(\rho^{\prime\prime\otimes n}))\geq$ (42)
${\frac{1}{n}}D_{\bf{C}}(\Lambda\circ{\cal M}(\rho^{\otimes n}\otimes\rho^{\prime\otimes n})||\Lambda\circ{\cal M}(\rho^{\prime\prime\otimes n}))\geq$ (43)
@@
where $\Lambda$ is on optimal one-way protocol realizing $R^{\rightarrow}(\rho,\rho^{\prime})$, and $\tilde{\gamma}^{Rn}$ is a state close to a private state by $\epsilon$. The state $\tilde{\sigma}_{n}$ is an arbitrary state. In the above, we have used monotonicity of the locally measured regularized relative entropy of entanglement under joint application of a CPTP map.
@@
$\frac{1}{n}D({\cal S}(\tilde{\gamma}^{Rn})||{\cal S}(\tilde{\sigma}_{n}))\geq$ (46)
$\frac{1}{n}\inf_{\sigma\in\mbox{Cone}(SEP,{\cal S}(\tilde{\sigma}_{n}))}D({\cal S}(\tilde{\gamma}^{Rn})||\sigma)\geq$ (47)
$\frac{1}{n}E_{R}^{\mbox{Cone}}(\psi_{+}^{Rn})-{1\over n}O(\epsilon)\geq$ (48)
$\frac{1}{n}[n\frac{\alpha-1}{\alpha}(R-\frac{1}{n}\tilde{E}_{\alpha}((\rho^{\prime\prime}\otimes\rho^{\prime\prime})^{\otimes n})]+\frac{1}{n}O(\epsilon+h(\epsilon)).$ (49)
In equation (46), we use joint monotonicity of the relative entropy under a privacy squeezing map ${\cal S}$, i.e., the inverse operation to twisting and tracing out the shielding system. In Eq. (47) we relax the relative entropy distance to the infimum over a convex set - the cone of states obtained by admixing any separable state to the state ${\cal S}(\tilde{\sigma}_{n})$. The latter state forms the apex of this set. Since the relative entropy distance from a bounded convex set that contains a maximally mixed state is asymptotically continuous Winter (2016), the subsequent inequality follows. In the last inequality, we use Lemma 2 and the fact that $\tilde{E}_{\alpha}$ is subadditive (see Eq. $5.26$ of Wilde et al. (2017)). ∎
@@
$E_{D}^{A\rightarrow B}(\rho_{AB})\geq D_{A}^{\infty}(\rho||{\hat{\rho}}),$ (50)
we have immediate corollary stated below.
@@
For a key-correlated states $\rho,\rho^{\prime}$ given in Eq. (12) and its attacked version $\hat{\rho},\hat{\rho}^{\prime}$ there is
@@
Proof.
We identify $A$ and $B$ from Eq. (50) with ${\bf C}\equiv C_{1}C_{2}$, and $AB$ respectively, to bound $D^{\infty}_{\bf C}(\rho\otimes\rho^{\prime})$ by $E_{D}^{\rightarrow}(\rho\otimes\rho^{\prime})$ from above. We further use Theorem 40. ∎
We can introduce some simplifications by considering the extremal value of $\alpha$ and behavior of the entanglement monotones involved in the above corollary. This leads us to a relaxed bound for a single state:
@@
It follows straightforwardly from the Corollary 51. Indeed, consider first the case $\rho=\rho^{\prime}$, and note that $\tilde{E}_{\alpha}$ is subadditive (see Eq. $5.26$ of Wilde et al. (2017)) and $E_{D}^{\rightarrow}$ is additive by definition. This implies
$R^{\rightarrow}(\rho,\rho)\leq 2({\alpha\over\alpha-1})E_{D}^{\rightarrow}(\rho)+2\tilde{E}_{\alpha}(\hat{\rho}).$ (53)
Relaxing the above inequality by taking the limit of $\alpha\rightarrow\infty$ we arrive at the claim. ∎
A special case of a key correlated state is a private bit (a private state with key part of local dimension $d_{k}=2$). When the private state is taken at random, we can then upper bound the term $E_{max}$ to obtain the bound which is dependent only on the one-way distillable entanglement and a constant factor. We thus arrive at the second main result of this section.
@@
In what follows, for the ease of notation, we will refer to $d_{s}$, i.e. the dimension of the shielding system of the key correlated state, as to $d$. We will relax the upper bound given in Eq. (52) provided in the Corollary 52. Let us recall first that $E_{max}(\rho):=\inf_{\sigma\in SEP}\inf\{\lambda\in{\mathrm{R}}:\rho\leq 2^{\lambda}\sigma\}$. As a further upper bound one can use the norm $||.||_{\infty}$. This is because $E_{max}(\rho)\leq\log_{2}\left(d^{2}||\rho||_{\infty}\right)$. The latter inequality can be easily seen from the fact that the maximally mixed state $\mathds{1}\over d^{2}$ is separable, and it majorizes every state. We then set $\sigma=\frac{\mathds{1}}{d^{2}}$ in the definition of $E_{max}$.If we have then an upper bound on the maximal eigenvalue of $\rho$, denoted as $\kappa_{max}$ the value $\lambda=\log_{2}(d^{2}\kappa_{max}))$ satisfies $\rho\leq 2^{\lambda}\mathds{1}/d^{2}$, and hence $E_{max}\leq\log_{2}(d^{2}\kappa_{max})$. It is known Marčenko and Pastur (1967), that the maximal eigenvalue of a random state $\rho_{rand}$ of dimension $d^{2}$ is upper bounded by
@@
where $\rho_{i}$ are also random states. Hence its maximal eigenvalue is upper bounded by ${1\over 2d^{2}}$. We have then that $E_{max}\leq\log_{2}(d^{2}\times({2\over d^{2}}))=1$, hence the assertion follows. ∎
@@
The bound given in Eq. (54) is loose as it has a non-vanishing factor of $2$. However, it is valid for any dimension of the shielding system of the state, and this factor is independent of the dimension. It implies that no matter how large is the shielding system of a key correlated state, generically, it brings in at most 2 bits of the repeated key above the (twice) one-way distillable entanglement.
It turns out, however, (see Corollary 3) that there is a tighter bound, which can be obtained not from the upper bounding via relative entropy but squashed entanglement, Christandl and Müller-Hermes (2017); Tucci (2002). In the next section, we will show that the (half of) the mutual information of a random private state does not exceed the minimal value, which is key by more than $\frac{1}{4\ln 2}$. The bound will then follow from a trivial inequality $R^{\rightarrow}(\rho,\rho)\leq R(\rho,\rho)\leq K_{D}(\rho)$ Bäuml et al. (2015).

### 09 Mar 2026 v4

@@ IV Relaxed bound on one-way private key repeaters
We follow the proof of Lemma $7$ of Horodecki et al. (2009a). Let $\sigma_{n}\in\mbox{Cone}(SEP,\tilde{\sigma}_{n})$. Using the concavity of the logarithm function, we first note that
$D(\psi_{+}^{Rn}||\sigma_{n})=\underbrace{\operatorname{Tr}\psi_{+}^{Rn}\log\psi_{+}^{Rn}}_{=0}-\operatorname{Tr}\psi_{+}^{Rn}\log\sigma_{n}\geq$
$\qquad\qquad-\log\operatorname{Tr}\psi_{+}^{Rn}\sigma_{n}$ (40)
Since $\sigma_{n}\in\mbox{Cone}(SEP,\tilde{\sigma}_{n})$, it can be expressed as $\sigma_{n}=p\tilde{\sigma}_{n}+(1-p)\sigma_{SEP}$ for some $p\in[0,1]$ and some $\sigma_{SEP}\in SEP$. Thus, we can rewrite the above equation as follows
$D(\psi_{+}^{Rn}||\sigma_{n})\geq-\log\operatorname{Tr}\psi_{+}^{Rn}\left(p\tilde{\sigma}_{n}+(1-p)\sigma_{SEP}\right)=$ (41)
$-\log\left(p\operatorname{Tr}\psi_{+}^{Rn}\tilde{\sigma}_{n}+(1-p)\operatorname{Tr}\psi_{+}^{Rn}\sigma_{SEP}\right)\geq$ (42)
$\qquad-\log\left(\max\{\operatorname{Tr}\psi_{+}^{Rn}\tilde{\sigma}_{n},\operatorname{Tr}\psi_{+}^{Rn}\sigma_{SEP}\}\right)$ (43)
Lemma 7 of Horodecki et al. (2009a) states that for a separable state $\sigma_{SEP}$ there is $\operatorname{Tr}\psi_{+}^{Rn}\sigma_{SEP}\leq 2^{-Rn}$. Note that since $Rn\geq n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}\tilde{E}_{\alpha}(\sigma^{\otimes n}))$, then $\operatorname{Tr}\psi_{+}^{Rn}\sigma_{SEP}\leq 2^{-n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}\tilde{E}_{\alpha}(\sigma^{\otimes n}))}$. Moreover lemma 1 states that $\operatorname{Tr}\psi_{+}^{Rn}\tilde{\sigma}_{n}\leq 2^{-n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}\tilde{E}_{\alpha}(\sigma^{\otimes n}))}$. Thus, we can combine all of that to obtain
$D(\psi_{+}^{Rn}||\sigma_{n})\geq-\log\left(\max\{\operatorname{Tr}\psi_{+}^{Rn}\tilde{\sigma}_{n},\operatorname{Tr}\psi_{+}^{Rn}\sigma_{SEP}\}\right)\geq$ (44)
$-\log 2^{-n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}\tilde{E}_{\alpha}(\sigma^{\otimes n}))}=n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}\tilde{E}_{\alpha}(\sigma^{\otimes n})),$ (45)
which is true for every $\sigma_{n}\in\mbox{Cone}(SEP,\tilde{\sigma}_{n})$. Thus, taking the infimum on the left-hand side completes the proof.
We are ready to state the main theorem. It shows that the one-way repeater rate is upper bounded by the locally measured regularized relative entropy of entanglement (increased by a factor $\frac{\alpha}{\alpha-1}$) in addition to the sandwich relative entropy of entanglement. In what follows, we will use notation $\rho\equiv\rho_{AC_{1}}$ and $\rho^{\prime}\equiv\rho^{\prime}_{C_{2}B}$, as well as $\rho^{\prime\prime}\equiv\rho^{\prime\prime}_{ABC_{1}C_{2}}$ and $C_{1}C_{2}\equiv{\bf C}$.
@@
For any states $\rho_{AC_{1}},\rho^{\prime}_{BC_{2}},\rho^{\prime\prime}_{ABC_{1}C_{2}}$ and a real parameter $\alpha\in(1,\infty)$, there is
$R^{\rightarrow}(\rho,\rho^{\prime})\leq\frac{\alpha}{\alpha-1}D_{\bf{C}}^{\infty}(\rho\otimes\rho^{\prime}||\rho^{\prime\prime})+\tilde{E}_{\alpha}(\rho^{\prime\prime}).$ (46)
@@
Our proof closely follows Christandl and Ferrara (2017) (Corollary 25), with suitable modifications. We first note that there is the following chain of (in)equalities, as it is noted in Christandl and Ferrara (2017):
@@
${\frac{1}{n}}D({\cal M}_{\bf{C}}(\rho^{\otimes n}\otimes\rho^{\prime\otimes n})||{\cal M}_{\bf{C}}(\rho^{\prime\prime\otimes n}))\geq$ (48)
${\frac{1}{n}}D(\Lambda\circ{\cal M}_{\bf{C}}(\rho^{\otimes n}\otimes\rho^{\prime\otimes n})||\Lambda\circ{\cal M}_{\bf{C}}(\rho^{\prime\prime\otimes n}))\geq$ (49)
@@
where $\Lambda$ is an optimal one-way protocol realizing $R^{\rightarrow}(\rho,\rho^{\prime})$, and $\tilde{\gamma}^{Rn}$ is a state close to some private state $\gamma^{Rn}=(\tau|\psi_{+}\rangle\langle\psi_{+}|\otimes\rho_{A^{\prime}B^{\prime}}\tau^{\dagger})^{Rn}$ by $\epsilon$. The state $\tilde{\sigma}_{n}$ is equal to $\operatorname{Tr}_{C}\mathcal{L}(\rho^{\prime\prime\otimes n})$, where $\mathcal{L}$ is an arbitrary operation from $LOCC({\bf C}\rightarrow(A\leftrightarrow B))$. In the above, we have used the monotonicity of the locally measured, regularized relative entropy of entanglement under the joint application of a CPTP map.
@@
$\frac{1}{n}D(\tilde{\gamma}^{Rn}||\tilde{\sigma}_{n})\geq\frac{1}{n}D({\cal S}(\tilde{\gamma}^{Rn})||{\cal S}(\tilde{\sigma}_{n}))\geq$ (52)
$\frac{1}{n}\inf_{\sigma\in\mbox{Cone}(SEP,{\cal S}(\tilde{\sigma}_{n}))}D({\cal S}(\tilde{\gamma}^{Rn})||\sigma)=\frac{1}{n}E_{R}^{\mbox{Cone}}({\cal S}(\tilde{\gamma}^{Rn}))$ (53)
In equation (52), we use joint monotonicity of the relative entropy under a privacy squeezing map ${\cal S}(\cdot)=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau^{\dagger}(\cdot)\tau$, i.e., the inverse operation to twisting $\tau$ composed with tracing out the shielding system. The subsequent inequality follows from the relaxation of entropy distance to the infimum over a convex set - the cone of states obtained by admixing any separable state to the state ${\cal S}(\tilde{\sigma}_{n})$. The latter state forms the apex of this set.
We continue lowerbounding ${\frac{1}{n}}D_{\bf{C}}(\rho^{\otimes n}\otimes\rho^{\prime\otimes n}||\rho^{\prime\prime\otimes n})$ as follows
$\frac{1}{n}E_{R}^{\mbox{Cone}}({\cal S}(\tilde{\gamma}^{Rn}))\geq$ (55)
$\frac{1}{n}\left[E_{R}^{\mbox{Cone}}(\psi_{+}^{Rn})-O(\epsilon)-h\left(\frac{\varepsilon}{1+\varepsilon}\right)\right]\geq$ (56)
$\frac{1}{n}\left[n\frac{\alpha-1}{\alpha}\left(R-\frac{1}{n}\tilde{E}_{\alpha}(\rho^{\prime\prime\otimes n})\right)\right]$- (57)
$\qquad\qquad{1\over n}\left[O(\epsilon)+h\left(\frac{\varepsilon}{1+\varepsilon}\right)\right]\geq$ (58)
$\frac{\alpha-1}{\alpha}\left(R-\tilde{E}_{\alpha}(\rho^{\prime\prime})\right)-{1\over n}\left[O(\epsilon)+h\left(\frac{\varepsilon}{1+\varepsilon}\right)\right].$ (59)
Inequality (55) follows from the fact that the relative entropy distance from a bounded convex set containing a maximally mixed state is asymptotically continuous (see Lemma 7 Winter (2016)). Next, since ${\cal S}(\tilde{\sigma}_{n})=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau^{\dagger}\operatorname{Tr}_{C}\mathcal{L}(\rho^{\prime\prime\otimes n})\tau$, we use Lemma 2. The last inequality follows from subadditivity of $\tilde{E}_{\alpha}$ (see Eq. $5.26$ of Wilde et al. (2017)). By rearranging the obtained inequality, we get
$\frac{\alpha}{\alpha-1}{\frac{1}{n}}D_{\bf{C}}(\rho^{\otimes n}\otimes\rho^{\prime\otimes n}||\rho^{\prime\prime\otimes n})+\tilde{E}_{\alpha}(\rho^{\prime\prime})\geq$ (60)
$\qquad\qquad\qquad R-{1\over n}\frac{\alpha}{\alpha-1}\left[O(\epsilon)+h\left(\frac{\varepsilon}{1+\varepsilon}\right)\right].$ (61)
Taking the limit $n\to\infty$, we end the proof.
@@
[…]
we have an immediate corollary stated below.
@@
For a key-correlated states $\rho_{AC_{1}},\rho^{\prime}_{BC_{2}}$ given in Eq. (13) and its attacked version $\hat{\rho}_{AC_{1}},\hat{\rho}^{\prime}_{BC_{2}}$ there is
$R^{\rightarrow}(\rho,\rho^{\prime})\leq\frac{\alpha}{\alpha-1}E_{D}^{\rightarrow}(\rho\otimes\rho^{\prime})+\min\{\tilde{E}_{\alpha}(\hat{\rho}),\tilde{E}_{\alpha}(\hat{\rho}^{\prime})\}.$ (63)
Proof.
We identify $A$ and $B$ from Eq. (62) with ${\bf C}\equiv C_{1}C_{2}$, and $AB$ respectively, to bound $D^{\infty}_{\bf C}(\rho\otimes\rho^{\prime}\|\hat{\rho}\otimes\hat{\rho}^{\prime})$ by $E_{D}^{\rightarrow}(\rho\otimes\rho^{\prime})$ from above. We further use Theorem 46 to obtain the following inequality
@@
Observe now that in all of the previous proofs we define $\tilde{E}_{\alpha}$ as an infimum with respect to states separable in $AC_{1}C_{2}:B$ cut. Thus we can bound $\tilde{E}_{\alpha}(\hat{\rho}_{AC_{1}}\otimes\hat{\rho}^{\prime}_{BC_{2}})$ as follows
$\tilde{E}_{\alpha}(\hat{\rho}_{AC_{1}}\otimes\hat{\rho}^{\prime}_{BC_{2}})\leq$ (65)
$\qquad\tilde{D}_{\alpha}(\hat{\rho}_{AC_{1}}\otimes\hat{\rho}^{\prime}_{BC_{2}}\|\hat{\rho}_{AC_{1}}\otimes\sigma_{BC_{2}})$ (66)
[…]
where $\sigma_{BC_{2}}$ is arbitrary separable state. Since it holds true for any choice of separable state $\sigma_{BC_{2}}$, we can write
$\tilde{E}_{\alpha}(\hat{\rho}_{AC_{1}}\otimes\hat{\rho}^{\prime}_{BC_{2}})\leq$ (68)
$\qquad\inf_{\sigma_{BC_{2}}\in\mathrm{SEP}}\tilde{D}_{\alpha}(\hat{\rho}^{\prime}_{BC_{2}}\|\sigma_{BC_{2}})=\tilde{E}_{\alpha}(\hat{\rho}^{\prime}_{BC_{2}}),$ (69)
which implies that
$R^{\rightarrow}(\rho,\rho^{\prime})\leq\frac{\alpha}{\alpha-1}E_{D}^{\rightarrow}(\rho\otimes\rho^{\prime})+\tilde{E}_{\alpha}(\hat{\rho}^{\prime}).$ (70)
On the other hand, notice that our considerations remain true if we define $\tilde{E}_{\alpha}$ as an infimum with respect to states separable in $A:C_{1}C_{2}B$ cut. Thus, by similar arguments, we can show that
[…]
By combining equations (70) and (71) we get
$R^{\rightarrow}(\rho,\rho^{\prime})\leq\frac{\alpha}{\alpha-1}E_{D}^{\rightarrow}(\rho\otimes\rho^{\prime})+\min\{\tilde{E}_{\alpha}(\hat{\rho}),\tilde{E}_{\alpha}(\hat{\rho}^{\prime})\},$ (72)
which finishes the proof.
We can introduce some simplifications by considering the extremal value of $\alpha$ and the behavior of the entanglement monotones involved in the above corollary. This leads us to a relaxed bound for a single state.
@@
It follows straightforwardly from the Corollary 1. Indeed, consider first the case $\rho=\rho^{\prime}$, and note that $\tilde{E}_{\alpha}$ is subadditive (see Eq. $5.26$ of Wilde et al. (2017)) and $E_{D}^{\rightarrow}$ is extensive by definition. This implies
[…]
Relaxing the above inequality by taking the limit of $\alpha\rightarrow\infty$, we arrive at the claim. ∎
A special case of a key correlated state is a private bit (a private state with a key part of local dimension $d_{k}=2$). When the private state is taken at random, we can then upper bound the term $E_{max}$ to obtain the bound which is dependent only on the one-way distillable entanglement and a constant factor. We thus arrive at the second main result of this section.
@@
[…]
@@
where $\rho_{i}$ are also random states. Hence, its maximal eigenvalue is upper bounded by $1\over 2d^{2}$}. We have then that $E_{max}\leq\log_{2}(d^{2}\times({2\over d^{2}}))=1$, hence the assertion follows. ∎
@@
The bound given in Eq. (75) is loose as it has a non-vanishing factor of $1$. However, it is valid for any dimension of the shielding system of the state, and this factor is independent of the dimension. It implies that no matter how large is the shielding system of a key correlated state, generically, it brings in at most $1$ bits of the repeated key above the (twice) one-way distillable entanglement.
Besides that, there is an independent bound (see Corollary 3), which can be obtained not from the upper bounding via relative entropy but squashed entanglement, Christandl and Müller-Hermes (2017); Tucci (2002). In the next section, we will show that the (half of) the mutual information of a random private state does not exceed the minimal value, which is key by more than $\frac{1}{4\ln 2}$. The bound will then follow from a trivial inequality $R^{\rightarrow}(\rho,\rho)\leq R(\rho,\rho)\leq K_{D}(\rho)$ Bäuml et al. (2015).